OpenStudy (anonymous):

Sally’s father baked 8 cookies. Sally and 4 friends (5 people total) decide to share the cookies equally. How much will each person get? A. 2 2 cookies B. 5/8 of a cookie C. 1 3/5 cookies D. 1 5/8 cookies

hero (hero):

So basically, compute 8 ÷ 5

hero (hero):

I can show you a simple way to compute it

hero (hero):

\[\frac{8}{5} = \frac{5+3}{5} = \frac{5}{5} + \frac{3}{5} = 1 \frac{3}{5}\]

hero (hero):

Any questions?